OpenStudy (anonymous):

find the slope of the line that goes through the given points. (-5,-5) and (-2,4)

OpenStudy (anonymous):

slope formula is \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

i dont understand so how am i to find the slope

OpenStudy (anonymous):

when you plug in the numbers it will look like this\[\frac{ 4-(-5) }{ -2-(-5)}\]

OpenStudy (anonymous):

now subtract

OpenStudy (anonymous):

-1,3

OpenStudy (anonymous):

4-(-5)=9 -2-(-5)=3

OpenStudy (anonymous):

so then the slope would be 9,3 xbeing 9 and y being 3

OpenStudy (anonymous):

so you have the fraction \[\frac{ 9 }{ 3 }\] just reduce and that's urslope

OpenStudy (anonymous):

so 1/3

OpenStudy (anonymous):

3/1

OpenStudy (anonymous):

right sorry thank you

OpenStudy (anonymous):

No problem :)